OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side. One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side. @Mathematics

OpenStudy (anonymous):

5n = n + 16

OpenStudy (anonymous):

n = 4

OpenStudy (anonymous):

so one side is equal to 20

OpenStudy (anonymous):

Wait how did you get n = 4?

OpenStudy (anonymous):

\[5n - n = 16 => 4n = 16 => n = \frac{16}{4} = 4\]